Will rate!!

Some parts of California are particularly earthquake prone. Suppose that in one metropolitan area, the chance a homeowner is insured against an earthquake is 0.30, A sample of four homeowners are to be selected at random. Suppose X is a random variable that is modeled by a binomial distribution which describes the number of homeowners out of the four that have earthquake insurance. (a) Find the probability mass function of X. (Round your answers to four decimal places.) p(x)
(b) Which of the following is a graph of the probability mass function (pmf)? Probablity Probability 0.4 0.3 0.2 0.1 0.3 Number with 0. 1 earthquake 2 3insurance Number with earthquake 2 3 4insurance Probability Probability 0.4 0,3 0.2 0.1 0.4 0.3 0.2 0.1 Number with earthquake insurance Number with earthquake insurance
(c) Which of the following is a graph of the cumulative distribution function (cdr)? Percentage Percentage 100 80 60 40 20 100 80 60 40 20 Number with earthquake insurance Number with earthquake S4 Insurance 0 12 3 4 Percentage Percentage 100 80 60 40 20 100 80 60 40 20 Number with Number with earthquake 3Insurance earthquake 12 3 4 insurance
(d) What is the most likely value for X? (e) What is the probability that at most 1 of the four selected have earthquake insurance? (Round your answer to four decimal places.) (1 What is the probability that at least two of the four selected have earthquake insurance? (Round your answer to four decimal places.) (9) What is the expected value and standard deviation of X? (Round your answer to two decimal places.)
Answer #1

from binomial distribution P(X=x)=4Cx(0.3)x(0.7)4-x

a)

x 0 1 2 3 4
P(x) 0.2401 0.4116 0.2646 0.0756 0.0081

b)top left curve is right

c) bottom right curve is right

d)

most likely value =1

e)

P(X<=1) =P(X=0)+P(X=1)=0.6517

f)

P(X>=2)=1-P(X<=1)=1-0.6517=0.3483

g)

E(X)=np=4*0.3=1.2

SD(X)=sqrt(np(1-p))=0.92